OpenStudy (anonymous):

Jesse rents a moving van for $75 and must pay $2 per mile. The following week, Alex rents the same van, is charged $80 for the rental and $1.50 per mile. If they each paid the same amount and drove the same number of miles, how far did they each travel?

OpenStudy (anonymous):

75+2x=80+1.5x 2x-1.5x=80-75 .5x=5 x=5/.5 x=10 miles.

OpenStudy (anonymous):

Hey, do you have anymore like this??

OpenStudy (anonymous):

Yeah, I have the one i just posted, and one more. Then I have I think 5 more...